# LVLM-COUNT: Enhancing the Counting Ability of Large Vision-Language Models

### v1

@@ Abstract
Counting is a fundamental skill for various visual tasks in real-life applications, requiring both object recognition and robust counting capabilities. Despite their advanced visual perception, large vision-language models (LVLMs) struggle with counting tasks, especially when the number of objects exceeds those commonly encountered during training. We enhance LVLMs’ counting abilities using a divide-and-conquer approach, breaking counting problems into sub-counting tasks. Unlike prior methods, which do not generalize well to counting datasets on which they have not been trained, our method performs well on new datasets without any additional training or fine-tuning. We demonstrate that our approach enhances counting capabilities across various datasets and benchmarks.
@@ 1 Introduction
Counting is a key cognitive task with broad applications in industry, healthcare, and environmental monitoring (De Almeida et al., 2015; Guerrero-Gómez-Olmedo et al., 2015; Paul Cohen et al., 2017; Lempitsky and Zisserman, 2010). It improves manufacturing, inventory, and quality control, ensures safety in medical settings, and helps manage resources in environmental efforts (Wang and Wang, 2011; Zen et al., 2012; Arteta et al., 2016). Recent advancements in prompt-based models enable counting of unlimited object varieties without visual exemplars. Although current trained, and text-prompt-based, counting models by Dai et al. (2024); Amini-Naieni et al. (2024) perform well on the datasets they are trained on, they face the following challenges. First, they require fine-tuning on new datasets. Second, since the concepts in the counting datasets are limited, these models do not generalize well to counting questions that involve complex reasoning. There are also training-free models by Shi et al. (2024), but overall they have weaker performance compared to trained models. On the other hand, large vision-language models (LVLMs), such as GPT-4o (Achiam et al., 2023), which also do not need dataset specific training, show very good performance in counting low numbers of objects, usually less than 20, across most datasets. However, their performance deteriorates for larger numbers of objects, regardless of the dataset.
We enhance the accuracy of LVLMs to count objects in images by leveraging their reasoning power within a divide-and-conquer method. The LVLMs allow us to handle diverse objects and complex counting questions, while our divide-and-conquer method alleviates challenges associated with counting large numbers of objects. Inspired by prior work on the rapid and accurate estimation of small quantities by Chattopadhyay et al. (2017), we divide an image into sub-images, and prompt the LVLM to count the objects of interest in each sub-image. The counts from each sub-image are then aggregated to make the final prediction. Our proposed workflow is illustrated in Figure 1.
@@
Figure 1: Illustration of our proposed pipeline. First, the area of interest, such as a set of zebras, is detected by extracting the object of interest ($E$) from a prompt question ($Q$) using a large language model (LLM) which is the same as LVLM in our work. Then, $E$ and the image are provided as input to an area detection model, such as the one by Liu et al. (2023). Second, any objects corresponding to $E$ are segmented. Third, in the object-aware division step, we use the segmentation masks to divide the areas of objects of interest without cutting through the objects. Finally, the number of objects of interest in each sub-image is computed using an LVLM, and the results are aggregated.
@@
Initially, in our pipeline, the object of interest is extracted from the input question using an LLM. The area containing the object of interest is detected in the image by a grounding model, such as Liu et al. (2023), and then cropped. The cropping step is important since it removes irrelevant context from the image. Secondly, using an object detection model by Liu et al. (2023), and a segmentation model by Kirillov et al. (2023), the segmentation masks of the objects of interest are created. Thirdly, we use a mechanism that divides the image into multiple sub-images without cutting through the objects of interest. We call this mechanism object-aware division. The division positions can be either set manually or determined automatically using unsupervised and non-parametric methods. In the latter case, we automatically select the division positions based on object masks, then we treat the object-aware division as a path-finding problem, avoiding objects as obstacles. A black-white image is built by converting all the masks into black and the rest of the image into white pixels. The binary image is converted into a graph where only white pixels are connected as nodes. Using the $A^{*}$ algorithm (Russell and Norvig, 2016), a path is found from one end to the other end of the image, ensuring objects remain intact. Finally, using an LVLM as a counting tool, the objects of interest in the sub-images are counted and aggregated. Our contributions are summarized below:
@@
    We propose LVLM-Count that leverages the strengths of large vision-language models (LVLMs) in visual perception for counting. Our method is a prompt-based counting approach that, in addition to simple counting problems, can handle complex cases as well. Our method does not require any additional training or fine-tuning on any dataset. LVLM-Count outperforms existing state-of-the-art models across various datasets and benchmarks.
@@
    We propose a solution for object-aware division. Accurate division is crucial, as parts of cut objects can lead to over-counting (see Figure 2(a)). To the best of our knowledge, this is the first method to divide images without cutting through objects of interest.
@@
As a minor contribution, we create a new benchmark to address two drawbacks of existing benchmarks. Prior datasets either feature simple counting tasks, e.g., counting “strawberries”, or include complex questions with small numbers of objects. To address both issues, we develop a challenging benchmark for counting emoji icons. The subtle variations within emoji classes make this benchmark uniquely difficult. Since none of the models have been exposed to it, this benchmark serves as a fair test of the performance of counting methods for various attributes and complex concepts.
@@ 2 Related Work
Early counting models, referred to as class-specific, targeted counting problems for certain categories (Arteta et al., 2016; Babu Sam et al., 2022; Mundhenk et al., 2016; Xie et al., 2018), such as cars, people, or cells. Later, with the emergence of stronger vision models and large-scale datasets, class-agnostic methods were proposed that could count objects from a wide variety of categories. However, most existing class-agnostic, or open-world, models require visual exemplars of the target objects (Đukić et al., 2023; Gong et al., 2022; Lin et al., 2022; Liu et al., 2022; Lu et al., 2019; Nguyen et al., 2022; Ranjan et al., 2021; Shi et al., 2022; Yang et al., 2021; You et al., 2023). The concept of divide and conquer has also been used in early work(Xiong et al., 2019; Chattopadhyay et al., 2017). However, this early work requires training dedicated models to utilize this concept.
@@
Text-based counting-specific trained models. With the advent of vision-language foundation models such as CLIP and GroundingDINO, text-based open-world methods have been proposed. Leveraging the rich textual and visual feature extraction capabilities of foundation models, obtained through web-scale training, the text-based counting methods by Amini-Naieni et al. (2023); Dai et al. (2024); Kang et al. (2024); Amini-Naieni et al. (2024) have started to demonstrate comparable or superior accuracy. GroundingREC by Dai et al. (2024) is an open-world model built on top of GroundingDINO (Liu et al., 2023), and introduces an additional task called referring expression counting. Concurrently, Amini-Naieni et al. (2024) proposed a method that also builds on GroundingDINO but adds an extra image-text fusion module in the input, enabling the model to accept text and/or visual exemplars to determine the target.
Models without counting-specific training. Shi et al. (2024) introduce TFOC, a counting model that does not require any counting-specific training. Instead, they cast the counting problem as a prompt-based segmentation task, using SAM (Kirillov et al., 2023) to obtain segmentation maps that determine the output number. Another group of models that do not need further training to count are LVLMs. State-of-the-art models such as GPT-4o (Achiam et al., 2023), Gemini 1.5 Pro (Reid et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024) show strong performance in counting small numbers of objects, although their performance degrades with larger numbers of objects. In comparison to prior work, our method (LVLM-Count) is an open-world, text-based counting approach that does not require any counting-specific training. Although we use pre-trained GroundingDINO and SAM in our method, we differ from prior work by not treating the counting task as a segmentation or detection problem. Instead, we use LVLMs as a counting tool.
@@ 3 LVLM-Count
Our proposed method aims to answer counting questions by dividing an image into sub-images while avoiding cuts through objects of interest. LVLM-Count consists of four key stages. First, in the “Area Detection” stage, we localize areas containing relevant objects. Second, in the “Target Segmentation” stage, we identify and segment these local areas. Third, in the “Object-aware Division” stage, we divide the localized areas into sub-images without cutting through the segmented objects. Finally, the LVLM counts the target objects in each sub-image and aggregates the results. Figure 1 illustrates the workflow of our method, which we will detail in the following subsections. Note that in this section, we illustrate each stage using example images. These images are are used solely for illustration purposes. LVLM-Count may not outperform other methods on these examples.
@@ 3.1 Area Detection
In this part of the pipeline, we assume that we are given a counting question $Q$ along with an image. The question $Q$ contains an expression $E$ that specifies a set of objects of interest. The expression $E$ distinguishes these objects from objects of other categories or the same category but with different attributes present in the image. By employing an LLM, the expression $E$ is extracted from $Q$. For example, let $Q$ be “How many people are in the boat?”. $Q$ is given to an LLM, which is prompted to return the expression $E$ “people in the boat”, referring to the objects we want to count. After $E$ is extracted, it is given as input to GroundingDINO along with the image. The output of GroundingDINO is a set of bounding boxes that have relevance to $E$ beyond a certain threshold. These bounding boxes often overlap and typically contain repeated objects. Thus, all the overlapping output bounding boxes are merged. After merging, a set of non-overlapping areas of interest may remain. We consider the non-overlapping areas as “detected areas”, which are then cropped to be passed to the next stage. Note that the area detection stage is important as it extracts the area with the most relevant context for the counting question. This process is illustrated in Figure 3.
Figure 3: Illustration of the area detection step of LVLM-Count. For this image, $Q$ is set to “How many people are in the boat”. The LLM that is used in this step returns an $E$ which is “people in the boat”. $E$ and the original image are given as input to GroundingDINO, which returns some bounding boxes (left and upper right images) that are merged to form the final detected area.
@@ 3.2 Target Segmentation
(a) Original image
(b) GroundingDINO output
(c) SAM output
@@
Figure 4: Illustration of the target segmentation step of LVLM-Count. The goal is to produce all the instance masks for $E$ set to “brown egg”. Figure 4(a), together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure 4(b). Figure 4(b) is then given as input to SAM, which produces the output shown in Figure 4(c).
@@
How to determine the bounding boxes. To determine the bounding boxes, we use GroundingDINO and set the bounding box probability threshold to a low value to avoid missing any objects. The bounding boxes alone cannot help with the object-aware division of the cropped images due to their rigid structure, which includes redundant areas in the vicinity of the object and, in the worst case, overlaps with other bounding boxes. Our goal is to precisely locate the pixels of an object of interest.
How to determine the segmentation masks. We use a pre-trained segmentation model, i.e., SAM (Kirillov et al., 2023), for the segmentation task, which accepts bounding boxes as prompts. It produces a mask covering the most prominent object within a bounding box. We run non-maximum suppression on the masks produced by SAM to remove the masks corresponding to less certain bounding boxes that have large overlaps with other masks. The masks produced for each cropped image are then passed to the next stage. For GroundingDINO we set the bounding box threshold to a very low value to avoid missing any objects of interest, especially if the objects belong to a rare category that GroundingDINO has encountered less frequently during pre-training. As shown in Figure 4, a low threshold value can lead to false positives, making the number of masks unreliable for the counting problem. Nonetheless, the false positive masks have no negative effect on the object-aware division in the next stage. The only consequence is that, in addition to the target masks, the division paths will not cut through the false positive masks either. In other words, extra masks are of no concern as long as all the target masks are identified.
@@ 3.3 Object-aware Division
How to determine the starting and ending points of the paths. We consider the first sub-step as a hyper-parameter, which can be set in various ways depending on the type of dataset. We utilize two approaches in our experiments. The first approach is unsupervised and non-parametric, where we use a clustering method to obtain the start and end points of the paths. We describe the clustering approach in more detail below. The second approach is to simply use a pre-determined number of equidistant points. This is helpful when we know a priori that objects are uniformly distributed across the image. In this case, we fix $k$ equidistant vertical and horizontal coordinates in the image as the points of the dividing paths for the entire dataset. In our experiments, we specify which approach we use for each dataset. Nonetheless, we will illustrate in the experimental results that even if we do not treat this as a hyperparameter and always use the unsupervised and non-parametric approach, LVLM-Count is quite effective.
Figure 5: Illustration of the unsupervised and non-parametric method to obtain the division points ($P^{1}_{s},P^{1}_{e}$), and ($P^{2}_{s},P^{2}_{e}$,). The object of interest is “horse”. Firstly, the target masks are produced. Then, a few pixels are sampled (shown as points inside the segmented objects) from the pixels composing each mask. The samples are projected onto the $x$-axis. The projected points are clustered using mean-shift clustering. The point in the middle of two consecutive clusters is considered a vertical division point. The straight vertical lines are obtained using our path-finding method.
@@
Description of the unsuperivsed and non-parametric approach. A few pixels are sampled from each of the masks. To determine the points of the vertical division paths, the samples taken from the masks are projected onto the $x$-axis. The projected points are automatically clustered using a non-parametric mean-shift algorithm. Once the clusters are identified, the point between the point with the highest $x$ value in one cluster and the point with the lowest $x$ value in the subsequent cluster is considered the $x$-coordinate of a vertical division path. Using this technique, we obtain the appropriate coordinates for the vertical paths, as well as the number of paths. For example, if there is only one cluster, no vertical division is required, and if there are two clusters, one vertical path will divide the image into two parts. We illustrate this approach in Figure 5. Note that we chose the vertical axis purely for convenience in illustration. The same process can be applied along the $y$-axis to obtain the $y$-coordinates of the horizontal division paths.
@@
How to draw the paths. In effect, knowing the $x$-coordinate of a vertical path means that the coordinates of its endpoints are known. In particular, assuming height $h$ for an image crop, we consider $P_{s}=(x,0)$ and $P_{e}=(x,h)$ as the start and end points, respectively. In an ideal case where there are no masks in the path of a straight line connecting the two points, this line will be drawn by connecting all the pixels on the straight path. However, there are potential masks that can be considered obstacles blocking the path. In other words, beginning from $P_{s}$, the line needs to go around these obstacles to reach $P_{e}$. Consequently, we treat this as a 2-dimensional path-finding problem. To solve the problem, we build a 2D binary map, $I_{B}$, where the pixels covered by the masks are turned into black, indicating them as obstacles, and all the other pixels are turned into white, showing they are open for passage. This binary image $I_{B}$ is mapped into a graph $G$, where each white pixel is a node, and it is connected to all of its white neighboring pixels. We use the $A^{*}(G,P_{s},P_{e},g)$ search algorithm to find a path that connects $P_{s}$ to $P_{e}$, where the heuristic $g$ is set to be Manhattan distance. The output of $A^{*}$ is a set of connected pixels that go around the obstacles and connect $P_{s}$ to $P_{e}$, creating an object-aware division path, as shown in Figure 6. The path-finding algorithm is run for all division coordinates. Finally, we draw the image contours based on these division paths and take the area surrounded by each contour as a resulting sub-image.
@@
Figure 6: Illustration of object-aware division. The masks for “donut” are obtained and turned into a black-and-white image. A dividing path is found by connecting $P_{s}$ to $P_{e}$ using the $A^{*}$ search algorithm in a graph that corresponds to the binary image, where the only nodes in the graph are white pixels, which are fully connected. The path is mapped back to the pixel domain.

### v3

@@ Abstract
Counting is a fundamental operation for various real-world visual tasks, requiring both object recognition and robust counting capabilities. Despite their advanced visual perception, large vision-language models (LVLMs) are known to struggle with counting tasks. In this work, we evaluate the performance of several recent LVLMs on visual counting tasks across multiple counting and vision datasets. We observe that while their performance may be less prone to error for small numbers of objects, they exhibit significant weaknesses as the number of objects increases. To alleviate this issue, we propose a simple yet effective baseline method that enhances LVLMs’ counting ability for large numbers of objects using a divide-and-conquer approach. Our method decomposes counting problems into sub-tasks. Moreover, it incorporates a mechanism to prevent objects from being split during division, which could otherwise lead to repetitive counting—a common issue in a naive divide-and-conquer implementation. We demonstrate the effectiveness of this approach across various datasets and benchmarks, establishing it as a valuable reference for evaluating future solutions.
@@ 1 Introduction
Counting is a key cognitive task with broad applications in industry, healthcare, and environmental monitoring (De Almeida et al., 2015; Guerrero-Gómez-Olmedo et al., 2015; Paul Cohen et al., 2017; Lempitsky and Zisserman, 2010). It improves manufacturing, inventory, and quality control, ensures safety in medical settings, and helps manage resources in environmental efforts (Wang and Wang, 2011; Zen et al., 2012; Arteta et al., 2016).
@@
Figure 1: Illustration of our proposed pipeline. First, an expression ($E$) describing the area of interest is extracted from the prompted question ($Q$), such as “brown eggs”. The expression is extracted using a large language model (LLM) which is the same as LVLM in our work. Then, $E$ and the image are provided as input to a grounding model, such as the one by Liu et al. (2023) to detect the area of interest. Second, any objects corresponding to $E$ are segmented. Third, in the object-aware division step, we use the segmentation masks to divide the detected area of interest without cutting through the objects of interest. Finally, the number of objects of interest in each sub-image is computed using an LVLM, and the results are aggregated.
Recent advancements in the field of large vision-language models (LVLMs), such as GPT-4o (Achiam et al., 2023), coupled with the massive scale of web-scraped data used to train them, have resulted in unprecedented zero-shot capabilities across diverse tasks. However, evaluations of LVLMs’ capabilities have revealed weaknesses in numerical reasoning Yin et al. (2023); Xu et al. (2024); Yang et al. (2023).
In this work, we focus specifically on the visual counting task, one of the most fundamental aspects of numerical reasoning. We observe that although LVLMs perform well in counting small numbers of objects—typically fewer than 20—their accuracy deteriorates with larger quantities. Inspired by prior work on the rapid and accurate estimation of small quantities by Chattopadhyay et al. (2017) we propose a simple yet effective baseline method to alleviate this issue. Leveraging a divide-and-conquer approach, we engineer a simple pipeline that divides an image into sub-images, and prompts the LVLM to count the objects of interest in each sub-image. The counts from the sub-images are then aggregated to make the final prediction. A key feature of our pipeline is a mechanism that prevents objects of interest from being split by the dividing lines, which could otherwise lead to double-counting. To have this mechanism two existing pre-trained detection and segmentation models are employed in the pipeline. The workflow is illustrated in Figure 1.
@@
Initially, in our pipeline, the category name of the object of interest is extracted from the input question using an LLM (We use the LVLM as an LLM for this step). The area containing the objects of interest is detected in the image by a grounding model, such as Liu et al. (2023), and then cropped. The cropping step removes irrelevant context from the image. Secondly, using an object detection model by Liu et al. (2023), and a segmentation model by Kirillov et al. (2023), the segmentation masks of the objects of interest are created. Thirdly, we use a mechanism that divides the image into multiple sub-images without cutting through the objects of interest. We call this mechanism object-aware division. The division positions are determined automatically using an unsupervised and non-parametric method based on object masks. Then we treat the object-aware division as a path-finding problem, avoiding the segmented objects as obstacles. A black-white image is built by converting all the masks into black and the rest of the image into white pixels. The binary image is converted into a graph where only white pixels are connected as nodes. Using the $A^{*}$ algorithm (Russell and Norvig, 2016), a path is found from one end to the other end of the image, ensuring objects remain intact. Finally, using an LVLM as a counting tool, the objects of interest in the sub-images are counted and aggregated. Our contributions are summarized below:
@@
    We evaluate the counting performance of several recent Large Vision-Language Models (LVLMs) on multiple counting and vision datasets. We propose a simple yet effective baseline method, LVLM-Count, which enhances the counting performance of LVLMs without requiring additional training. Similar to standard counting with LVLMs, our method is a prompt-based approach that retains their zero-shot capabilities while addressing their difficulties in handling large numbers. Through experiments, we demonstrate that LVLM-Count improves LVLMs’ counting performance across the evaluated datasets.
@@
    We propose a solution for object-aware division. Accurate division is crucial, as parts of cut objects can lead to over-counting (see Figure 2(a)). The proposed solution divides images without cutting through objects of interest specified by an arbitrary prompt.
@@
As a minor contribution, we create a new benchmark to address some of the drawbacks of existing datasets. Prior datasets feature simple counting tasks, e.g., counting “strawberries”, and lack intra-class complexity. To address these issues, we develop a challenging benchmark for counting emoji icons. The subtle variations within emoji classes make this benchmark uniquely difficult.
@@ 2 Related Work
Early counting models, referred to as class-specific, targeted counting problems for certain categories (Arteta et al., 2016; Babu Sam et al., 2022; Mundhenk et al., 2016; Xie et al., 2018), such as cars, people, or cells. Later, with the emergence of stronger vision models and large-scale datasets, class-agnostic methods were proposed that could count objects from a wide variety of categories. However, most existing class-agnostic, or open-world, models require visual exemplars of the target objects (Đukić et al., 2023; Gong et al., 2022; Lin et al., 2022; Liu et al., 2022; Lu et al., 2019; Nguyen et al., 2022; Ranjan et al., 2021; Shi et al., 2022; Yang et al., 2021; You et al., 2023).
@@
Text-based counting models. With the advent of vision-language foundation models such as CLIP and GroundingDINO, text-based open-world models have been proposed that are trained specifically for counting. Leveraging the rich textual and visual feature extraction capabilities of foundation models, obtained through web-scale training, the text-based counting models by Amini-Naieni et al. (2023); Dai et al. (2024); Kang et al. (2024); Amini-Naieni et al. (2024) have started to demonstrate comparable or superior accuracy. In addition, Shi et al. (2024) introduce TFOC, a counting model that does not require any counting-specific training. Instead, they cast the counting problem as a prompt-based segmentation task, using SAM (Kirillov et al., 2023) to obtain segmentation masks that determine the output number.
Leveraging the concept of divide and conquer for counting. The concept of divide and conquer has been used in early work (Chattopadhyay et al., 2017; Xiong et al., 2019; Stahl et al., 2018). Chattopadhyay et al. (2017) use an image-level divide and conquer approach and train a convolutional neural network (CNN) that can count objects from a predetermined and limited set of categories in sub-images. Xiong et al. (2019) propose applying the divide step on the convolutional feature map instead of the input image to avoid repeatedly computing convolutional features for each sub-image, thereby improving efficiency. However, the CNN in their work is only capable of counting a single object category. Similar to Chattopadhyay et al. (2017), Stahl et al. (2018) also employ image-level division and train a CNN to count objects from a predetermined set of categories. Nonetheless, their method does not require local image annotations for training.
Assessment of LVLMs’ counting performance. Several prior works have explored the visual counting capabilities of LVLMs as part of broader evaluations, underscoring the difficulties these models encounter in counting tasks Yin et al. (2023); Xu et al. (2024); Yang et al. (2023). However, these studies have not focused on developing solutions to address these challenges. In this work, we conduct a comprehensive quantitative assessment of LVLMs’ counting performance across diverse visual counting benchmarks. More importantly, we propose a simple yet effective baseline method to improve their ability to count large numbers. Our baseline method is not tailored to any specific model and can be used in a plug-and-play manner with different LVLMs. Regarding the categories outlined earlier, our method, LVLM-Count, is an open-world, prompt-based counting approach that requires no additional training. To the best of our knowledge, we are the first to propose a divide-and-conquer strategy that avoids splitting and potentially double-counting objects of interest specified by an arbitrary text prompt.
@@ 3 LVLM-Count
Our proposed method aims to answer counting questions by dividing an image into sub-images while avoiding cuts through objects of interest. LVLM-Count consists of four key stages. First, in the “Area Detection” stage, we localize areas containing relevant objects. Second, in the “Target Segmentation” stage, we identify and segment the objects of interest. Third, in the “Object-aware Division” stage, we divide the localized areas into sub-images without cutting through the segmented objects. Finally, the LVLM counts the target objects in each sub-image and aggregates the results. Figure 1 illustrates the workflow of our method, which we detail in the following subsections.
@@ 3.1 Area Detection
In this part of the pipeline, we assume that we are given a counting question $Q$ along with an image. The question $Q$ contains an expression $E$ that specifies a set of objects of interest. The expression $E$ distinguishes these objects from objects of other categories or the same category but with different attributes present in the image. By employing an LLM, the expression $E$ is extracted from $Q$. For example, let $Q$ be “How many brown eggs are in the image?”. $Q$ is given to an LLM, which is prompted to return the expression $E$, “brown eggs”, referring to the objects we want to count. After $E$ is extracted, it is provided as input to GroundingDINO along with the image. The output of GroundingDINO may be a single bounding box or a set of bounding boxes that have relevance to $E$ beyond a certain threshold. These bounding boxes often overlap and typically contain repeated objects. Thus, all the overlapping output bounding boxes are merged. After merging, a set of non-overlapping areas of interest may remain. We consider the non-overlapping areas as “detected areas”, which are then cropped to be passed to the next stage. This process is illustrated in Figure 3.
Figure 3: Illustration of the area detection stage of LVLM-Count. For this image, $Q$ is set to “How many brown eggs are in the image”. The LLM that is used in this step returns an $E$ which is “brown eggs”. $E$ and the original image are given as input to GroundingDINO, which returns a bounding box. If the grounding model returns multiple bounding boxes, they are merged to form the final detected area.
@@ 3.2 Target Segmentation
(a) GroundingDINO output
(b) SAM output
@@
Figure 4: Illustration of the target segmentation step of LVLM-Count. The goal is to produce all the instance masks for $E$ set to “brown egg”. The cropped detected area from Figure 3, together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure 4(a). Figure 4(a) is then given as input to SAM, which produces the output shown in Figure 4(b).
@@
How to determine the bounding boxes. To determine the bounding boxes, we use GroundingDINO and set the bounding box probability threshold to a low value to avoid missing any object of interest. The bounding boxes alone cannot help with the object-aware division of the detected areas due to their rigid structure, which includes redundant areas in the vicinity of the object and, in the worst case, overlaps with other bounding boxes. Our goal is to precisely locate the pixels of an object of interest.
How to determine the segmentation masks. We employ a pre-trained segmentation model, specifically SAM (Kirillov et al., 2023), for the segmentation task. This model accepts bounding boxes as prompts and generates masks covering the most prominent objects within these boxes. However, we do not use SAM’s output masks directly, as in crowded scenes or cases with multiple occlusions, the masks often overlap, making it difficult to identify reliable division paths. To address this issue, we apply several post-processing steps. First, we perform non-maximum suppression on SAM-generated masks to eliminate those with significant overlaps corresponding to less certain bounding boxes. Additionally, we apply an erosion function to the segmentations, ensuring adjacent masks maintain a minimum separation of two pixels. For a detailed analysis and visual examples of how these post-processing operations enhance robustness in crowded scenes and occlusion cases, see Appendix C. The final processed masks for each cropped area are then passed to the next stage of our pipeline.
Robustness to the Accuracy of Area Detection and Target Segmentation Stages.
Although we employ GroundingDINO—a state-of-the-art model for grounding and detection—it lacks the flexibility and robustness of LVLMs when generalizing to unseen data and complex concepts. Consequently, we minimizes dependence on the accuracy of the detection model used for area detection and target segmentation. Rather than prioritizing accuracy, we emphasize: i) not missing any region containing objects of interest during area detection, and ii) not missing segmentation masks for any objects of interest during target segmentation. These objectives are easily achieved by setting GroundingDINO’s detection threshold to a very low value in both stages.
A low threshold may lead to false positives. In area detection, this could result in regions that contain no objects of interest. However, since counting is performed by the LVLM in the final stage, these regions will be ignored. For target segmentation, false positive masks might occur, but the consequence is that some irrelevant objects will be protected from being cut by the division lines, just as the objects of interest are. Furthermore, as we will demonstrate in subsequent sections, all masks are removed from the sub-images after division, ensuring no noise propagates from this stage to the LVLM-based counting stage.
We demonstrate the effectiveness and robustness of our strategy in the early pipeline stages by showing that LVLM-Count significantly improves LVLMs’ performance on one of the most challenging counting datasets—the Penguin Dataset Penguin Research (2016), as detailed in Section 4.2. This dataset is particularly difficult due to frequent occlusion, camouflage, and complex backgrounds Arteta et al. (2016). Furthermore, in our ablation studies, we test a variant of our pipeline that removes the detection model entirely. Instead, SAM is configured to generate segmentation masks for all objects in the scene. As shown in Table 5, our method still enhances counting performance in this configuration, underscoring its robustness and minimal dependence on the detection model during the initial pipeline stages.
@@ 3.3 Object-aware Division
How to determine the starting and ending points of the paths. We utilize an unsupervised and non-parametric approach, to obtain the start and end points of the paths. A few pixels are sampled from each of the masks. To determine the location of the division paths on the $x$-axis, the samples taken from the masks are projected onto the $x$-axis. The projected points are automatically clustered using a non-parametric mean-shift algorithm. Once the clusters are identified, the point between the point with the highest $x$ value in one cluster and the point with the lowest $x$ value in the subsequent cluster is considered the $x$-coordinate of a division path. In effect, knowing the $x$-coordinate of a vertical path means that the coordinates of its endpoints are known. In particular, assuming height $h$ for an image crop, we consider $P_{s}=(x,0)$ and $P_{e}=(x,h)$ as the start and end points, respectively. Note that using this technique, we obtain the appropriate coordinates for the division paths, as well as the number of paths. For example, if there is only one cluster along the $x$-axis, no division is required, and if there are two clusters, one vertical path will divide the image into two parts. We illustrate this approach in Figure 5.
Figure 5: Illustration of the unsupervised and non-parametric method to obtain the division points ($P^{1}_{s},P^{1}_{e}$), and ($P^{2}_{s},P^{2}_{e}$,). A few pixels are sampled (shown as points inside the segmented objects) from the pixels composing target masks. The samples are projected onto the $x$-axis. The projected points are clustered using mean-shift clustering. The point in the middle of two consecutive clusters is considered a vertical division point. Blue lines are solely for illustration
@@
How to draw the paths. Previous step obtains the endpoints of the division paths. Assume $P_{s}=(x,0)$ and $P_{e}=(x,h)$ are the start and end points of a vertical division path, respectively. In an ideal case where there are no masks in the path of a straight line connecting the two points, this path will be drawn by connecting all the pixels on the straight line. However, there are potential masks that can be considered obstacles blocking the path. In other words, beginning from $P_{s}$, the line needs to go around these obstacles to reach $P_{e}$. Consequently, we treat this as a 2-dimensional path-finding problem. To solve the problem, we build a 2D binary map, $I_{B}$, where the pixels covered by the masks are turned into black, indicating them as obstacles, and all the other pixels are turned into white, showing they are open for passage. This binary image $I_{B}$ is mapped into a graph $G$, where each white pixel is a node, and it is connected to all of its white neighboring pixels. We use the $A^{*}(G,P_{s},P_{e},g)$ search algorithm to find a path that connects $P_{s}$ to $P_{e}$, where the heuristic $g$ is set to be Manhattan distance. The output of $A^{*}$ is a set of connected pixels that go around the obstacles and connect $P_{s}$ to $P_{e}$, creating an object-aware division path, as shown in Figure 6. The path-finding algorithm is run for all division coordinates. Finally, we draw the image contours based on these division paths and take the area surrounded by each contour as a resulting sub-image.
@@
Figure 6: Illustration of object-aware division. The masks are turned into a black-and-white image. A dividing path is found by connecting $P_{s}$ to $P_{e}$ using the $A^{*}$ search algorithm in a graph that corresponds to the binary image. The only nodes in the graph are white pixels of the black-and-white image, which are connected to all other white pixels in their $3\times 3$ neighborhood. The nodes and edges on the obtained paths have been colored red.